OpenStudy (anonymous):

What is the sum of one over six plus two over three in simplest form? Answer one over six one over three two over three five over six

OpenStudy (lgbasallote):

\[\huge \frac 16 + \frac 23 \implies \frac 16 + \frac 46 \implies \frac{1+4}{6} \implies \frac 56\] does that help?

OpenStudy (lgbasallote):

@twil you there?